OpenStudy (anonymous):

Need Help :) A peanut butter jar has a height of 7.2in. and a diameter of 4.1in. One cubic inch holds. 5oz of peanut butter. How many ounces will fit in the jar? Round to the nearest whole number.

OpenStudy (anonymous):

so the volume of a cylinder is \[\pi \times d \times h\] d is 3.14 and h is 7.2 \[\pi \times 7.2 \times 4.2\] which equals 92.739 cubic inches so because each 5 of those holds 1 oz you divide by 5\[92.379 / 5 = 18.54\] which is closest to 19 oz of peanut butter

OpenStudy (anonymous):

Thanks :)